Station ALE, Alert, NU, Canada

ALE operational status:


The station was powered down for building maintenance activities from day 2016,173 (June 21) to 2016,193.


The station was powered back up after the vault construction was completed on day 2015,238 (Aug. 26).


The station is down while the structure surrounding the vault is rebuilt.


The station operator re-centered the STS-2.5.  All channels from the STS-2.5 are seismic again.


The masses for the STS-2.5 seismometer have drifted out of normal operating range.  Instructions for re-centering the sensor have been sent to the station operator.


IDA engineers are onsite to upgrade the station to a Q330HR system.  Sensors include STS-1 and STS-2.5 seismometers, and a Setra micro-barometer.


IDA station ALE is down after the station operator temporarily decommissioned the station. There are plans to upgrade the station to a Quanterra Q330HR system early next year.


The GPS clock is unlocked as of day 2011,282 due to a software problem. IDA engineers are working to resolve this issue.


The GPS clock at the station was unlocked as of day 2011,049-12:58. The station operator reports cycling power and restoring normal operation of the GPS on day 2011,091.


The mass positions for the STS-2 were saturated from about day 2010,275-0800 to 2010,337-1429, when the station operator re-centered the masses for the STS-2.


The station was down from day 2010,221-16:44 to 2010,263-11:54 due to problems with the ISP. Data acquisition was restored after the station operator replaced the ISP.


The station operator reports the station is back on-line after reconfiguring the ISP. The station operator will send all available data so we may determine the extent of time the station may have been down.


The station operator reports problems with the new stealth station processor. The station is currently down and is not acquiring data. IDA engineers are working with the station operator to reconfigure the station computer.


Station operator reports receiving a new stealth station processor shipped from UCSD.


The station operator cycled power to the digitizer and reports the station is back online.


The station is down as of about day 2007,181-21:22 due to communication problems between the DAS and the ISP that are yet to be determined.


The GPS clock is in an unlocked state as of about day 160-01:01. Timing of data since this time is uncertain. The station operator has been requested to cycle power to the GPS clock. Data are continuous and data waveform quality remain unaffected.


The station operator reports that the CD drive for the ISP failed.


There are problems with the station’s GPS clock.


Data acquisition restored after comm reset by operator. DAS firmware upgraded to address recording of wdi data.


The VMV-10 and VMW-10 channels from the STS-2 seismometer are saturated because the sensor’s masses need to be recentered.


Telemetry problems persist. The station operator requested to reset comm (10/1).


Telemetry back after station operator cycled power to the iridium modem and router.


Telemetry down as of 2003:239-22:36 (Aug.27). Station operator requested to cycle power on transceiver and router.


The data acquisition system has been upgraded to a MK8 system. Additional sensors have been added including STS2 and FBA-23 seismometers and a Paroscientific microbarograph.


Station is operating normally.

Low gain broadband channels show problems with multiplexor from 2000.01.01. High gain broadband channels from 2000.02.25. To be fixed upon upgrade.